OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - x2 = -3x - 10?

OpenStudy (anonymous):

Move the x to the other side.

OpenStudy (anonymous):

Then you have a quadratic

OpenStudy (anonymous):

Solve for x by either factoring OR by the quadratic equation

OpenStudy (anonymous):

You should get x= -2, 5

OpenStudy (anonymous):

write in standard form: y=x^2-3x-10 then to find x-intercepts you need to solve 0=x^2-3x-10

OpenStudy (anonymous):

y=x^2-3x-10 y=(x-5)(x+2) x intercepts are x=5 and x=-2

OpenStudy (anonymous):

WOOT. Yep yep